# Low-cost Embedded Breathing Rate Determination Using 802.15.4z IR-UWB Hardware for Remote Healthcare

### v1

@@ 6.2 Generalization to new situations
Compared to prior work, the performance of the rule-based solutions are very close to the results reported in the papers from Farsaei et al. [13] and Numan et al. [20] for the “simple” situations 1 and 2. When considering more difficult situations, which are not considered in the works above, the algorithms start to fail, whereas the CNN still provides reasonable accuracies for these challenging, unseen conditions.
@@ 7.2 Algorithmic complexity of preprocessing pipeline
This section theoretically analyzes which parameters most profoundly impact the computation complexity of the preprocessing pipeline. First, the pipeline applies an FFT to each 30-second range bin. This FFT operation has a complexity of $O(n~{}\log~{}n)$, with $n$ the array length. Since an FFT is performed for each fast-time index, indicated by $m$, the total complexity becomes $O(mn\log~{}n)$. Next, a cropping operation reduces the window size from $(n,m)$ to $(k,r)$, where $k$ is the number of frequency bins and $r$ is the number of range bins after cropping. This slicing operation has a complexity of $O(kr)$. Finally, a normalization operation is applied by using an element-wise subtraction and division operation. Both these operations have a complexity of $O(k~{}r)$. All previous steps can be summarized as $O(mn~{}\log~{}n+3kr)$. Since $m$ and $r$ are radio configuration-dependent constants, this simplifies to $O(n\log~{}n+k)$. The dominant term, being $O(n\log~{}n)$, indicates an acceptable computational load suitable for constrained hardware.
@@ 7.5.1 Experimental measurement of the energy consumption
This section examines the energy consumption of the solution. The practical energy consumption was determined by deploying all components of the processing pipeline on the nRF52840 DK platform and measuring the current draw with a Keysight N6705B DC Power Analyzer. As the hardware was powered with a USB cable, the supply voltage was set to 5 V. Note that the currents mentioned here are averages, while the reported energy consumptions are precise area-under-the-curve calculations. Total energy consumption per 30-second window can be calculated using Equation 7.
$\begin{split}E_{30sec}=&\underbrace{E_{plat}+E_{TX}}_{TX}\\ &+\underbrace{E_{plat}+E_{RX}+E_{preproc}+E_{model}}_{RX}\end{split}$ (7)
This formula includes the energy consumption of the platform ($E_{plat}$), transmitter ($E_{TX}$), receiver ($E_{RX}$), preprocessing pipeline ($E_{preproc}$), and CNN model ($E_{model}$). Since the transmitter and receiver are separate devices, the platform’s energy must be counted twice. As all CIR s are collected at the receiver, processing will occur at the receiver side. The nRF52840 DK platform consumes $1.84~{}mA$ on average at $5~{}V$, resulting in an energy consumption of $E_{plat}=0.277~{}J$ over 30 seconds.
As shown in Section 7.1, the system works well at a sampling rate of 20 Hz. In this case, transmitter and receiver consume $0.0447~{}J$ and $0.0637~{}J$ per second, respectively. Over 30 seconds, this totals to $E_{TX}=1.342~{}J$ and $E_{RX}=1.912~{}J$. Preprocessing occurs once every 30 seconds, consuming $2.01~{}mA$ over $0.110~{}s$, which equals $E_{preproc}=0.001~{}J$. The CNN analysis uses $4.56~{}mA$ over $0.523~{}s$, consuming $0.012~{}J$. Quantizing the model reduces current draw to $4.17~{}mA$ and inference time to $0.199~{}s$, reducing energy consumption to $E_{model}=0.004~{}J$. Total energy consumption over 30 seconds is $E_{30sec}=3.813~{}J$. In other words, the system consumes $0.127~{}W$ on average. With an AAA battery providing $5400~{}J$, 1 battery can power the system for approximately 11.8 hours. Since each battery provides only $1.5~{}V$, at least 4 batteries are needed to power the $5~{}V$ system. Section 7.1 also states that the sampling rate can be lowered to 4 Hz in more simple situations. This lowers power consumption of the transmitter and receiver to $0.0436~{}J$ and $0.0547~{}J$ per second, respectively. Over 30 seconds, this becomes $E_{TX}=1.307~{}J$ and $E_{RX}=1.640~{}J$. The preprocessing step was also tested at this lowered sampling rate, but no measurable differences were observed. Recalculating with these numbers, show that the system consumes $0.117~{}W$ at a sampling rate of 4 Hz. This corresponds to 12.8 hours of constant monitoring with 4 AAA batteries.
@@ 7.5.2 Theoretical energy consumption model
(i) A first optimization is to deploy the transmitter and receiver on 1 platform. By doing this, only 1 $E_{plat}$ term remains in the energy calculation. To this end, different antenna designs should minimize self-interference between the transmitting and receiving antenna [24].
@@
Theoretical durations and current usages are presented in Table 4. Note that the transceiver operates at 3V, so all energy calculations were performed with this voltage in mind. The transmitter starts in the PLL state, using 18 mA for 0.020 ms. It then transmits the SHR, which is 256 symbols long, taking 0.270 ms and consuming 48 mA. Next, the PHR and PSDU are sent. For CIR collection, empty packets can be sent, but the PSDU will still contain the medium access control (MAC) header and reed solomon (RS) bits, taking 0.016 ms and consuming 40 mA. The transmitter then transitions into a deep sleep state for the rest of the 250 or 50-millisecond period, based on the configured sampling rate. In this case, 4 Hz and 20 Hz are shown. This state is very efficient, with a consumption of only 0.00026 mA. The receiver also starts in the PLL state and then skips the HUNT state. After this, the receiving device sequentially acquires the SHR, PHR, and PSDU, ending with another PLL, IDLE RC, and deep sleep state. When considering the system with a sampling rate of 20 Hz, 1 propagation sums to $0.0419~{}mJ$ for the transmitter and $0.0626~{}mJ$ for the receiver. This is executed 20 times per second over a period of 30 seconds, resulting in $E_{TX\_min}=0.025~{}J$ and $E_{RX\_min}=0.0375~{}J$. Substituting these theoretical minima in Equation 7, gives $E_{30\_sec\_min}=0.345~{}J$. Consequently, on average 0.0115 W is used, allowing the optimized system to run for 130.5 hours on four AAA batteries. A more realistic deployment option is a common battery pack of 20 000 mAh. This configuration could enable the system to work for 268.2 days without recharging. In less complex situations, where a sampling rate of 4 Hz is sufficient, $E_{TX\_min}$ and $E_{RX\_min}$ decrease to $0.005~{}J$ and $0.008~{}J$. Consequently, $E_{30\_sec\_min}$ becomes $0.295~{}J$, enabling continuous monitoring for approximately 313.8 days. This allows for BR monitoring in locations without power outlets, such as movable beds, furniture, and desks.
@@ 8 Conclusion
This research successfully developed and validated a non-intrusive system for accurate BR monitoring. The proposed solution has potential applications across diverse domains, including remote patient monitoring, stress detection in vehicles and ships, and presence detection in smart homes. A CNN model, trained and evaluated on a novel, open source dataset, demonstrated significantly superior performance compared to existing approaches. Robustness was extensively tested across challenging conditions. The system maintained accuracy at distances up to 4 meters, achieving an average L1 error of 2.41 BPM, which was further reduced to 1.29 BPM with calibration pretraining. Accurate results were also obtained with different body angles, showing a mean error of 0.86 BPM at 45 degrees. Moreover, the system demonstrated generalizability by accurately performing in unseen environments and with subjects in various postures. Deployment feasibility was shown using a low-cost, COTS radar setup and optimized processing, enabling operation on resource-constrained hardware. Practical analysis revealed 11.8 hours of continuous monitoring using 4 AAA batteries at a sampling rate of 20 Hz. Theoretical optimizations indicate the potential for significantly extended battery life, suggesting up to 130.5 hours of operation on AAA batteries, or 268.2 days with a common battery pack. For less complex situations, sampling rate can be lowered to 4 Hz. This decrease in sampling frequency extends the theoretical battery life to 313.8 days. Future research will address even more real-world complexities such as multi-person detection and non-static subjects and explore integration with communication modules for joint communication and sensing (JCAS) applications.

### v2

@@ 6.2 Generalization to new situations
Rule-based results in “simple” scenarios (setups 1 and 2) are comparable to prior work by Farsaei et al. [13] and Numan et al. [20]. However, in more challenging conditions not addressed by these prior studies, their performance degrades sharply, while the proposed CNN maintains reasonable accuracy. Validation folds for environments 1, 2, and 3 involved entirely unseen surroundings. The model’s strong performance in these cases demonstrates its robustness to environmental variation, a common challenge for many AI-based methods. Comparable evidence of generalizability is also absent in Zhou et al. [19].
@@ 7.2 Algorithmic complexity of preprocessing pipeline
This section theoretically analyzes which parameters most profoundly impact the computation complexity of the preprocessing pipeline. First, the pipeline applies an FFT to each 30-second range bin. This FFT operation has a complexity of $O(n~\log~n)$, with $n$ the array length [24]. Since an FFT is performed for each fast-time index, indicated by $m$, the total complexity becomes $O(mn\log~n)$. Next, a cropping operation reduces the window size from $(n,m)$ to $(k,r)$, where $k$ is the number of frequency bins and $r$ is the number of range bins after cropping. This slicing operation has a complexity of $O(kr)$. Finally, a normalization operation is applied by using an element-wise subtraction and division operation. Both these operations have a complexity of O(kr)O(k~r). All previous steps can be summarized as $O(mn~\log~n+3kr)$. Since $m$ and $r$ are radio configuration-dependent constants, this simplifies to $O(n\log~n+k)$. The dominant term, being $O(n\log~n)$, indicates an acceptable computational load suitable for constrained hardware.
@@ 7.5.1 Experimental measurement of the energy consumption
This section analyzes the energy consumption of the proposed solution. Practical measurements were obtained by deploying the complete processing pipeline on the nRF52840 DK platform and monitoring current draw with a Keysight N6705B DC Power Analyzer at a supply voltage of 5 V (USB-powered). Reported currents represent averages, while power values are derived from precise area-under-the-curve calculations. These measurements were combined into a full-system energy model (Equation 7), which accounts for platform ($P_{plat}$), transmitter ($P_{TX}$), and receiver ($P_{RX}$) consumption. Since preprocessing and model inference occur only once every 30 s, they are included as $E_{preproc}$ and $E_{model}$, respectively. As transmitter and receiver are separate devices, $P_{plat}$ is counted twice, with all processing executed on the receiver side.
Per-component measurements are summarized in Table 4. The nRF52840 DK consumes 1.84mA1.84~mA at 5 V. Preprocessing is executed once every 30 s, drawing 2.01mA2.01~mA for 0.110s0.110~s. CNN inference requires 4.56mA4.56~mA for 0.523s0.523~s, which is reduced to 4.17mA4.17~mA and 0.199s0.199~s after quantization, yielding substantial savings. At a sampling rate of 20 Hz, which is shown in Section 7.1 to provide robust operation under diverse conditions, the total energy consumption per 30 s is $E_{30sec}=3.813~J$, corresponding to an average power of 0.127W0.127~W. With a typical AAA battery providing 5400J5400~J, the system can operate for about 11.8 h. As each battery delivers only 1.5V1.5~V, at least four batteries are required to supply the 5 V system. When the sampling rate is reduced to 4 Hz for simpler scenarios, transmitter and receiver consumption decreases, lowering $E_{30sec}$ to 3.507J3.507~J (average power 0.117W0.117~W). Preprocessing at this rate showed no measurable difference. Under this configuration, four AAA batteries enable approximately 12.8 h of continuous monitoring.
$\begin{split}E_{\text{system}}&=\\ &\underbrace{(P_{\text{plat}}+P_{\text{TX}})\cdot t}_{\text{TX}}+\\ &\underbrace{(P_{\text{plat}}+P_{\text{RX}})\cdot t+(E_{\text{preproc}}+E_{\text{model}})\cdot\left\lfloor\tfrac{t}{30}\right\rfloor}_{\text{RX}}\end{split}$ (7)
V (V) P (W) T (s)
Practical
Platform 5 0.00923 continuous
Preprocessing 5 0.01016 0.110
Model 5 0.02287 0.523
Quantized model 5 0.02146 0.199
$TX_{20\_Hz}$ 5 0.04474 continuous
$RX_{20\_Hz}$ 5 0.06373 continuous
$TX_{4\_Hz}$ 5 0.04357 continuous
$RX_{4\_Hz}$ 5 0.05468 continuous
Theoretical
$TX_{20\_Hz}$ 3 0.000838 continuous
$RX_{20\_Hz}$ 3 0.001252 continuous
$TX_{4\_Hz}$ 3 0.000168 continuous
$RX_{4\_Hz}$ 3 0.000251 continuous
Table 4: Voltage and power consumption of components in the processing pipeline. Preprocessing and model stages consume power intermittently, while other modules draw continuous power. Practical results are based on measured energy usage from deployed components, whereas theoretical results represent lower bounds after applying the optimizations described in Section 7.5.
@@ 7.5.2 Theoretical energy consumption model
(i) A first optimization is to deploy the transmitter and receiver on 1 platform. By doing this, only 1 $P_{plat}$ term remains in the energy calculation. To this end, different antenna designs should minimize self-interference between the transmitting and receiving antenna [25].
@@
Theoretical durations and current usages are presented in Table 5. Note that the transceiver operates at 3V, so all energy calculations were performed with this voltage in mind. The transmitter starts in the PLL state, using 18 mA for 0.020 ms. It then transmits the SHR, which is 256 symbols long, taking 0.270 ms and consuming 48 mA. Next, the PHR and PSDU are sent. For CIR collection, empty packets can be sent, but the PSDU will still contain the medium access control (MAC) header and reed solomon (RS) bits, taking 0.016 ms and consuming 40 mA. The transmitter then transitions into a deep sleep state for the rest of the 250 or 50-millisecond period, based on the configured sampling rate. In this case, 4 Hz and 20 Hz are shown. This state is very efficient, with a consumption of only 0.00026 mA. The receiver also starts in the PLL state and then skips the HUNT state. After this, the receiving device sequentially acquires the SHR, PHR, and PSDU, ending with another PLL, IDLE RC, and deep sleep state. When considering the system with a sampling rate of 20 Hz, 1 propagation sums to 0.0419mJ0.0419~mJ for the transmitter and 0.0626mJ0.0626~mJ for the receiver. Multiplying this by 20 gives the power consumption per second as stated in Table 4. These new values give a new power consumption per 30 seconds: $E_{30\_sec\_min}=0.345~J$. Consequently, on average 0.0115 W is used, allowing the optimized system to run for 130.5 hours on four AAA batteries. A more realistic deployment option is a common battery pack of 20 000 mAh. This configuration could enable the system to work for 268.2 days without recharging. In less complex situations, where a sampling rate of 4 Hz is sufficient, $P_{TX\_min}$ and $P_{RX\_min}$ decrease even further. Consequently, $E_{30\_sec\_min}$ becomes 0.295J0.295~J, enabling continuous monitoring for approximately 313.8 days. This allows for BR monitoring in locations without power outlets, such as movable beds, furniture, and desks.
@@ 8 Conclusion
This research developed and validated a non-intrusive system for accurate BR monitoring with potential applications in remote patient monitoring, stress detection in vehicles or ships, and presence detection in smart homes. A CNN model, based on an architecture specifically adapted for IR-UWB CIR data and trained on a new open-source dataset, outperformed existing methods. Robustness was confirmed across distances up to 4 m (average L1 error 2.41 BPM, reduced to 1.29 BPM with calibration), body angles (0.86 BPM at 45°), unseen environments, and a variety of postures.
Complexity analysis verified the preprocessing pipeline is lightweight, and the CNN was effectively quantized with minimal accuracy loss, enabling deployment on a broader range of platforms. A novel analytical energy model was made and estimated 11.8 h of constant monitoring with four AAA batteries at 20 Hz, extendable to 130.5 h with theoretical optimizations, 268.2 days with a common battery pack, and 313.8 days at 4 Hz sampling rate. Results also show that sampling at 4 Hz is sufficient in less complex scenarios, whereas a higher rate of 20 Hz is recommended for analyzing more complex cases.
Future research can address multi-person detection, non-static subjects, and integration with joint communication and sensing (JCAS) modules. Explainable AI also represents a highly promising direction for future research, particularly in the context of healthcare applications.